OpenStudy (anonymous):

Write in terms of sine and cosine and simplify the expression. sinA- 2cosAsinA divide by sin^2(A)-cos^2(A)+cosA-1

myininaya (myininaya):

\[\frac{sinA-2cosAsinA}{\sin^2A-\cos^2A+cosA-1}\]

myininaya (myininaya):

\[\frac{sinA(1-2cosA)}{1-\cos^2A-\cos^2A+cosA-1}\]

myininaya (myininaya):

\[\frac{sinA(1-2cosA)}{-2\cos^2A+cosA}=\frac{sinA(1-2cosA)}{cosA(-2cosA+1)}\]

myininaya (myininaya):

\[=\frac{sinA}{cosA}\]

myininaya (myininaya):

any questions

OpenStudy (anonymous):

how did you get the 1 on the bottom?

myininaya (myininaya):

1-2cosA=-2cosA+1 they cancel addition is commutative

myininaya (myininaya):

examples: 3+2=2+3 3+(-2)=-2+3

OpenStudy (anonymous):

i see... okay.

myininaya (myininaya):

did you see i wrote sin^2A as 1-cos^2A

OpenStudy (anonymous):

is that because its an identity?

myininaya (myininaya):

yes!

OpenStudy (anonymous):

thank you

myininaya (myininaya):

np well you know sin^2a+cos^2a=1 right i just subtracted cos^2a on both sides to obtain sin^2a=1-cos^2a